a=6
(c) dx (x2+a2)y x2-a2 (d) 6xydy +(y2- (7 + a)x2 + axy )dx = 0
Answer #1

so you or age = *Cated you a great het om og det er mulig = a(a + 6) (n-6573 täccatou = a (476) I.f- espalda spatot a holda =2 oln bay da avtn. +nd GO 13-61 do- - ® Gay dy + (y²-13 n²t6xy) danso du may = 1337-6xy - y Y=0x = dy y oth doa Totado a = 13